### v1

# Association Scheme on Triples from the Unitary Group

Abstract
An association scheme on triples (AST) is a three-dimensional analogue of a classical association scheme. Similar to how a transitive group action produces a Schurian classical association scheme, a two-transitive group action produces an AST. This paper describes the relations and intersection numbers of ASTs from the finite unitary groups acting on the respective isotropic lines.
Keywords: association scheme on triples, unitary groups, isotropic lines
MSC 2020: 05E30
1 Introduction
A classical association scheme on a set $\Omega$ is a partition of $\Omega\times\Omega$ satisfying certain regularity properties. By providing a single framework with which to study several discrete structures, classical association schemes have proven useful in the study of combinatorial objects such as graphs, codes, and designs.
Introduced in [5] as a higher-dimensional analogue of a classical association scheme, an association scheme on triples (AST) on a set $\Omega$ is a partition of $\Omega\times\Omega\times\Omega$ satisfying certain symmetry requirements. Similar to how classical association schemes arise from transitive group actions, ASTs arise naturally from two-transitive group actions. The parameters of several families of such ASTs were obtained in [1].
However, the parameters of some families of such ASTs remain undetermined. Indeed, due to the more involved action of the finite unitary groups on isotropic lines, the parameters of the corresponding ASTs were left incomplete. In this paper, we continue the work in [1] by leveraging the matrix representation of the unitary groups to describe explicitly the relations and intersection numbers of the ASTs these groups produce. The parameters are expressed in the following main result.
Theorem 3.2. Let $q$ be a prime power, $U(3,q^{2})$ be the unitary group of $3\times 3$ matrices over $GF(q^{2})$, $\mathbb{P}C$ be the set of isotropic lines in the 3-dimensional vector space over $GF(q^{2})$, and $X$ be the AST from the $2$-transitive action of $U(3,q^{2})$ on $\mathbb{P}C$. Then the intersection numbers of $X$ are as listed in Table 1.
2 Preliminaries
2.1 Association Schemes on Triples
Following [5], we define an association scheme on triples as follows.
Definition 2.1.
Let $\Omega$ be a finite non-empty set with at least 3 elements. An association scheme on triples (AST) on $\Omega$ is a partition $\left\{R_{0},R_{1},R_{2},R_{3},\ldots,R_{m}\right\}$ of $\Omega\times\Omega\times\Omega$ such that $m\geq 4$ and the following conditions hold:
  1. i.
    For any $i\in\left\{0,1,2,3,\ldots,m\right\}$, there exists a non-negative integer $n_{i}$ such that for any distinct elements $x$ and $y$ in $\Omega$,
    $n_{i}=\left|\left\{\omega\in\Omega:(x,y,\omega)\in R_{i}\right\}\right|.$
  2. ii.
    For any $i,j,k,\ell\in\left\{0,1,2,3,\ldots,m\right\}$, there exists a non-negative integer $p_{ijk}^{\ell}$ such that for any $(x,y,z)\in R_{\ell}$,
    $p_{ijk}^{\ell}=\left|\left\{\omega\in\Omega:(\omega,y,z)\in R_{i},(x,\omega,z)\in R_{j},(x,y,\omega)\in R_{k}\right\}\right|.$
  3. iii.
    For any $i\in\left\{0,1,2,3,\ldots,m\right\}$ and any permutation $\sigma$ of the set $\left\{1,2,3\right\}$, there exists $j\in\left\{0,1,2,3,\ldots,m\right\}$ such that
    $R_{j}=\left\{(x_{\sigma(1)},x_{\sigma(2)},x_{\sigma(3)}):(x_{1},x_{2},x_{3})\in R_{i}\right\}.$
  4. iv.
    The trivial relations $R_{0}$, $R_{1}$, $R_{2}$, and $R_{3}$ are the following:
    $R_{0}$ $=\left\{(x,x,x):x\in\Omega\right\},$
    $R_{1}$ $=\left\{(y,x,x):x,y\in\Omega,x\neq y\right\},$
    $R_{2}$ $=\left\{(x,y,x):x,y\in\Omega,x\neq y\right\},$
    $R_{3}$ $=\left\{(x,x,y):x,y\in\Omega,x\neq y\right\}.$
The integer $n_{i}$ is called the third valency (or simply valency) of the ternary relation $R_{i}$. The integer $p_{ijk}^{\ell}$ is called the intersection number with respect to the relations $R_{i},R_{j},R_{k}$ and $R_{\ell}$. For ease of discussion, define the set
$\Gamma_{ijk}^{\ell}(x,y,z):=\left\{w\in\Omega:(w,y,z)\in R_{i},(x,w,z)\in R_{j},(x,y,w)\in R_{k}\right\},$
for any indices $i,j,k,\ell\in\{0,\ldots,m\}$ and any $(x,y,z)\in R_{\ell}$. By construction, $p_{ijk}^{\ell}=\left|\Gamma_{ijk}^{\ell}(x,y,z)\right|$. The following remark provides some readily computed parameters.
Remark 2.1 (Proposition $2.7$, [5]).
Suppose $X=\left\{R_{0},R_{1},R_{2},R_{3},\ldots R_{m}\right\}$ is an arbitrary AST on a set $\Omega$ for some $m\geq 4$. The following hold:
$n_{0}=n_{3}=0$
$n_{1}=n_{2}=1$
$p_{000}^{0}=1$
$p_{123}^{0}=\left|\Omega\right|-1$
$p_{011}^{1}=p_{202}^{2}=p_{330}^{3}=1$
$p_{132}^{1}=p_{321}^{2}=p_{213}^{3}=1$
$p_{ijk}^{1}=p_{ijk}^{2}=p_{ijk}^{3}=0$, where $i,j,k\geq 4$
$p_{3\ell 1}^{\ell}=p_{\ell 32}^{\ell}=p_{21\ell}^{\ell}=1$, where $\ell\geq 4$
Since several other intersection numbers are readily determined, the following intersection numbers are the ones of interest.
$p_{1jk}^{1}$, $p_{i2k}^{2}$, $p_{ij3}^{3}$, where $i,j,k\geq 4$
$p_{ijk}^{\ell}$, where $i,j,k,\ell\geq 4$
2.2 Algebra of Hypermatrices
To interpret the intersection numbers of an AST algebraically, we view ASTs as hypermatrix ternary algebras [5]. Let $X=\{R_{i}\}_{i=0}^{m}$ be an AST on a set $\Omega$. For each $i\in\{0,\ldots,m\}$, we define an $|\Omega|\times|\Omega|\times|\Omega|$ cubic hypermatrix $A_{i}$ whose entries are indexed by $\Omega$. The $(x,y,z)$-entry of $A_{i}$ is
$(A_{i})_{xyz}=\begin{cases}1,&\text{if }(x,y,z)\in R_{i},\text{ and }\\ 0,&\text{if }(x,y,z)\notin R_{i}.\end{cases}$
The linear space $\operatorname{Span}_{\mathbb{C}}(\{A_{i}\}_{i=0}^{m})$ of the adjacency hypermatrices is a complex vector space under scalar multiplication and entry-wise addition. To equip the span with a ternary algebra structure, define for each $A,B,C\in\operatorname{Span}_{\mathbb{C}}(\{A_{i}\}_{i=0}^{m})$ the product $ABC$ given by
$(ABC)_{xyz}=\displaystyle\sum_{w\in\Omega}{(A)_{wyz}(B)_{xwz}(C)_{xyw}}.$
The theorem below shows that the intersection numbers of $X$ are the structure constants of a ternary algebra.
Theorem 2.1 (Theorem $1.4$ and Corollary 2.8, [5]).
Let $X=\{R_{i}\}_{i=0}^{m}$ be an AST with corresponding adjacency hypermatrices $\{A_{i}\}_{i=0}^{m}$. Then $\operatorname{Span}_{\mathbb{C}}(\{A_{i}\}_{i=0}^{m})$ is a ternary algebra satisfying $A_{i}A_{j}A_{k}=\displaystyle\sum^{m}_{\ell=0}{p_{ijk}^{\ell}A_{\ell}},$ for any $i,j,k\in\left\{0,1,\ldots,m\right\}$. Moreover, $\operatorname{Span}_{\mathbb{C}}(\{A_{i}\}_{i=4}^{m})$ is a ternary subalgebra of $\operatorname{Span}_{\mathbb{C}}(\{A_{i}\}_{i=0}^{m})$.
In particular, Theorem 2.1 suggests that the most interesting intersection numbers are $p_{ijk}^{\ell}$ for $i,j,k,l\geq 4$.
2.3 ASTs from Two-Transitive Groups
Let $G$ be a group acting on a set $\Omega$. We say that $G$ is $\mathbf{2}$-transitive on $\Omega$ if for any $a,b,c,d\in\Omega$ with $a\neq b$ and $c\neq d$, there exists $x\in G$ such that $a^{x}=c$ and $b^{x}=d$. We write $a^{G}$ and $G_{a}$ for the orbit and the stabilizer of an element $a\in G$, respectively.
We state and outline a proof of the following theorem from [5], thereby including details on the valencies and intersection numbers.
Theorem 2.2 (Theorem $4.1$, [5]).
Let $G$ be a group acting $2$-transitively on a set $\Omega$, and let $\Omega^{3}$ denote $\Omega\times\Omega\times\Omega$. Define the action
$\Omega^{3}\times G$ $\longrightarrow\Omega^{3}$
$(a,b,c)$ $\xmapsto{\phantom{..}x\phantom{..}}{}(a^{x},b^{x},c^{x})$
of $G$ on $\Omega^{3}$. The set of orbits of this action on $\Omega^{3}$ is an AST on $\Omega$.
Outline of proof..
Suppose $\left\{R_{0},R_{1},R_{2},R_{3},\ldots,R_{m}\right\}$ is the set of distinct orbits of the action of $G$ on $\Omega^{3}$. Since Conditions (iiiiii) and ($iv$) of Definition 2.1 follow directly from the two-transitivity of $G$, we may assume that $R_{0},R_{1},R_{2}$, and $R_{3}$ are the trivial relations described in Definition 2.1. To verify Condition ($i$) of Definition 2.1, take any $a\neq b$ and $c\neq d$. Choose $x$ such that $(a^{x},b^{x})=(c,d)$. We then obtain the bijection
$\left\{e\in\Omega:(a,b,e)\in R_{i}\right\}$ $\longrightarrow\left\{f\in\Omega:(c,d,f)\in R_{i}\right\}$
$e$ $\longmapsto e^{x}.$
To verify condition ($ii$) of Definition 2.1, let $i,j,k,\ell\in\left\{0,1,\ldots,m\right\}$. Let $(a,b,c)$ and $(x,y,z)$ be in $R_{\ell}$. There exists $g\in G$ such that $(x,y,z)^{g}=(a,b,c)$. The map $w\mapsto w^{g}$ from $\Gamma_{ijk}^{\ell}(x,y,z)$ into $\Gamma_{ijk}^{\ell}(a,b,c)$ is a bijection. ∎
Remark 2.2 (Remark 2.10 [1], Lemma 4.2 [5]).
Let $G$ be a two-transitive group acting on a set $\Omega$, and $X$ be the AST from $G$ as constructed in Theorem 2.2. The verification of Condition $(ii)$ of Definition 2.1 in the above proof shows that the nontrivial relations of $X$ are in correspondence with the orbits of any two-point stabilizer. In fact, it also shows that the valencies are the sizes of these orbits.
2.4 AST from the Unitary Group
Let $q$ be a prime power, $F=GF(q^{2})$, and $\overline{c}=c^{q}$ for each $c\in F$. For any subfield $K$ of $F$, let $K^{\times}$ denote the multiplicative group of units of $K$. Further, let $F_{0}=\left\{r\in GF(q^{2}):r^{q}=r\right\}$ be the order $q$ subfield of $F$, and let $V=F\times F\times F$ be the usual 3-dimensional vector space over $F$. Define $B:V\times V\longrightarrow V$ by
$B(x,y)=x_{1}\overline{y_{2}}+x_{2}\overline{y_{1}}+x_{3}\overline{y_{3}},$
for each $x=(x_{1},x_{2},x_{3}),y=(y_{1},y_{2},y_{3})\in V$. Since $B$ is linear in the first variable and $\overline{B(x,y)}=B(y,x)$ for any $x,y\in V$, the mapping $B$ is a Hermitian form on $V$ [4]. In fact, $B$ is non-degenerate; i.e. if $x\in V$ satisfies $B(x,y)=0$ for all $y\in V$, then $x=0$.
Let $U(3,q^{2})$ be the group of all linear automorphisms $\sigma:V\rightarrow V$ such that
$B(\sigma(x),\sigma(y))=B(x,y).$
Then $U(3,q^{2})$ is called the unitary group of $V$. Letting $Z(U(3,q^{2}))$ denote the center of $U(3,q^{2})$, the quotient group $PGU(3,q^{2})=\frac{U(3,q^{2})}{Z(U(3,q^{2}))}$ is called the projective unitary group.
Let $x\in V$ be a nonzero vector and let $[x]$ be the one-dimensional subspace of $V$ spanned by $x$. If $B(x,x)=0$, then we say that $x$ is an isotropic vector and that $[x]$ is an isotropic line. The set $\mathbb{P}C$ of all isotropic lines is called the projective Hermitian quadratic cone. A subspace $W$ of $V$ is totally isotropic if every nonzero vector in $W$ is isotropic. The dimension of a maximal totally isotropic subspace is called the Witt index of $V$. It can be shown that $V$ has a Witt index of $1$, and that $V$ has a basis of vectors $u,v,w\in V$ satisfying $B(u,v)=1=B(w,w)$ and $B(u,u)=0=B(v,v)$. We fix and utilize such a basis $\{u,v,w\}$ throughout the article.
3 Parameters of the AST from the Unitary Group
The group $U(3,q^{2})$ acts two-transitively on $\mathbb{P}C$ via $\sigma[x]=[\sigma(x)]$ for $\sigma\in U(3,q^{2})$ [4]. It can be shown that the kernel of this action is $Z(U(3,q^{2}))$, so that $PGU(3,q^{2})$ acts faithfully and two-transitively on $\mathbb{P}C$. Let $\mathbb{P}C^{3}=\mathbb{P}C\times\mathbb{P}C\times\mathbb{P}C$ and define the following action
$\mathbb{P}C^{3}\times U(3,q^{2})$ $\longrightarrow\mathbb{P}C^{3}$
$([x],[y],[z])$ $\xmapsto{\phantom{..}\sigma\phantom{..}}{}([\sigma(x)],[\sigma(y)],[\sigma(z)])$
of $U(3,q^{2})$ on $\mathbb{P}C^{3}$. By Theorem 2.2, the set $X$ of orbits of this action forms an association scheme on triples AST on $\mathbb{P}C$. We compute the intersection numbers of this AST in this section. Through these computations, we recover the number of relations and the valencies of this AST, initially obtained in [1].
Theorem 3.1 (Table $1$, [1]).
Let $X$ be the AST from the $2$-transitive action of $U(3,q^{2})$. Then the number of relations of $X$ is $q+5$.
For our purposes, we require explicit forms for the nontrivial relations of this AST. To accomplish this, first define the sets
$\Omega_{0}:=\left\{[au+v]\in\mathbb{P}C:a\in F^{\times},a+\overline{a}=0\right\},$
and
$\Omega^{\prime}:=\left\{[bu+v+cw]\in\mathbb{P}C:b,c\in F,\;b+\overline{b}+c\overline{c}=0\neq b+\overline{b}\right\}.$
It follows that $\mathbb{P}C$ is the disjoint union of the sets $\left\{[u],[v]\right\}$, $\Omega_{0}$, and $\Omega^{\prime}$.
By Remark 2.2, there is a correspondence between the orbits of $U(3,q^{2})$ on $\mathbb{P}C^{3}$ and the orbits of $U(3,q^{2})_{[u],[v]}$ on $\mathbb{P}C\setminus\left\{[u],[v]\right\}$. Since it can be verified that $U(3,q^{2})_{[u],[v]}$ acts transitively on $\Omega_{0}$, it remains to consider the orbits of $U(3,q^{2})$ on $\Omega^{\prime}$. These are described in Proposition 3.1, affording us a workable description of the nontrivial relations.
Proposition 3.1.
The group $U(3,q^{2})_{[u],[v]}$ acts on $\Omega^{\prime}$, and this action has $q$ orbits.
Proof.
To obtain the number of orbits of $U(3,q^{2})_{[u],[v]}$ on $\Omega^{\prime}$, we begin by determining $|\Omega^{\prime}|$. Since the trace map $T:F\rightarrow F_{0}$ given by $T(x)=x+\overline{x}$ is an additive epimorphism, we have
$|\operatorname{Ker}{T}|=\frac{|F|}{|F_{0}|}=\frac{q^{2}}{q}=q.$ (1)
Moreover,
$F\setminus\operatorname{Ker}{T}=\{b\in F^{\times}:b+\overline{b}\neq 0\}.$
Since the norm map $N:F^{\times}\rightarrow F_{0}^{\times}$ with $N(r)=r\overline{r}$ is a multiplicative group epimorphism, we deduce that $|\operatorname{Ker}{N}|=q+1$. For every $b\in F\setminus\operatorname{Ker}{T}$, the set of all elements $c\in F$ such that $b+\overline{b}+c\overline{c}=0$ forms a coset of $\operatorname{Ker}{N}$. Hence, there are exactly $q+1$ such elements $c\in F$ satisfying $b+\overline{b}+c\overline{c}=0$.
Thus,
$|\Omega^{\prime}|=|F\setminus\operatorname{Ker}{T}|(q+1)=(q^{2}-q)(q+1)=q(q^{2}-1).$
Proceeding, observe that for any $bu+v+cw],[b_{1}u+v+c_{1}w]\in\Omega^{\prime}$, we have $bu+v+cw]\in[b_{1}u+v+c_{1}w]^{U(3,q^{2})_{[u],[v]}}$ if and only if $(F_{0}^{\times})b=(F_{0}^{\times})b_{1}$. It follows that any orbit contains exactly
$|F_{0}^{\times}|(q+1)=(q-1)(q+1)=q^{2}-1$ (2)
isotropic lines. Therefore,
$\frac{|\Omega^{\prime}|}{q^{2}-1}=\frac{q(q^{2}-1)}{q^{2}-1}=q$
is the number of orbits of $U(3,q^{2})_{[u],[v]}$ on $\Omega^{\prime}$. ∎
The proof of Proposition 3.1 implies that the orbits of $U(3,q^{2})_{[u],[v]}$ on $\Omega^{\prime}$ correspond to the cosets of $F_{0}^{\times}$ in $F^{\times}$. To enumerate the relations in $X$, let $\{b_{5},\ldots,b_{q+4}\}$ be a complete set of coset representatives for $F^{\times}/F_{0}^{\times}$. Further, fix for each $i\in\{5,\ldots,q+4\}$ an element $c_{i}\in F$ such that $b_{i}+\overline{b_{i}}+c_{i}\overline{c_{i}}=0$. Lastly, fix a nonzero $a\in F$ such that $a+\overline{a}=0$. The AST $X$ has two kinds of nontrivial relations. There is one of the first type, namely
$R_{4}=([u],[v],[au+v])^{U(3,q^{2})}.$ (3)
The remaining $q$ relations of $X$ are
$R_{i}=([u],[v],[b_{i}u+v+c_{i}w])^{U(3,q^{2})},$ (4)
where $i\in\{5,\ldots,q+4\}$. The valencies of the nontrivial relations are then obtained as the following consequence of Equations (1) and (2).
Corollary 3.1 (Table $1$, [1]).
Let $X$ be the AST from the $2$-transitive action of $U(3,q^{2})$. Then the nontrivial valencies are $n_{4}=q-1$ and $n_{i}=q^{2}-1$ for $i\geq 5$.
We may sharpen the description of the relations of the AST from $U(3,q^{2})$ when the underlying field has even characteristic.
Remark 3.1.
Let $q$ be a power of two, $X$ be the AST from the action of $U(3,q^{2})$, and $F$ be the field of cardinality $q^{2}$. Fix a primitive element $\alpha$ of $F$ and let $T:F\rightarrow F_{0}$ be given by $T(x)=x+\overline{x}$. Since $F_{0}^{\times}\subseteq\operatorname{Ker}T$, we may choose $a=\alpha^{q+1}$ and $b_{i}=\alpha^{(q+1)+i}$ for $i\in\{4,\ldots,q+4\}$ in Equations (3) and (4).
We proceed with the intersection numbers of the AST from $U(3,q^{2})$.
Theorem 3.2.
Let $q$ be a prime power, $U(3,q^{2})$ be the unitary group of $3\times 3$ matrices over $GF(q^{2})$, $\mathbb{P}C$ be the set of isotropic lines in the 3-dimensional vector space over $GF(q^{2})$, and $X$ be the AST from the $2$-transitive action of $U(3,q^{2})$ on $\mathbb{P}C$. For $5\leq i,j,k,\ell\leq q+4$, the intersection number $p_{ijk}^{\ell}$ of $X$ is as listed in Table 1.
1 $p_{144}^{1}=p_{424}^{2}=p_{443}^{3}=q-1$
2 $p_{14i}^{1}=p_{1i4}^{1}=p_{42i}^{2}=p_{i24}^{2}=p_{4i3}^{3}=p_{i43}^{3}=0$
[…]
For items 6 – 8, fix $([u],[v],[au+v])\in R_{4}$.
6 $p_{444}^{4}=q-2$
7 $p_{i44}^{4}=p_{4j4}^{4}=p_{44k}^{4}=p_{ij4}^{4}=p_{i4k}^{4}=p_{4jk}^{4}=0$
[…]
For items 9 – 13, fix $([u],[v],[b_{\ell}u+v+c_{\ell}w])\in R_{\ell}$.
9 $p_{i44}^{\ell}=p_{4j4}^{\ell}=p_{44k}^{\ell}=p_{444}^{\ell}=0$
[…]
13 $p_{ijk}^{\ell}=\left|\left\{([u],[v],[bu+v+cw])\in R_{k}:b_{i}bb_{\ell}^{-1}(\overline{b}+b_{\ell}+\overline{c}c_{\ell}),b_{j}(\overline{b}+b_{\ell}+\overline{c}c_{\ell})^{-1}\in F_{0}^{\times}\right\}\right|$
Table 1: Intersection numbers $p_{ijk}^{\ell}$ of the AST from $U(3,q^{2})$, where $5\leq i,j,k,\ell\leq q+4$.
Proof.
We only show that line 8 of Table 1 holds; the other lines of Table 1 can be shown by applying a similar approach. Let $([u],[v],[au+v])\in R_{4}$. Observe that $p_{ijk}^{4}\neq 0$ if and only if
$\Gamma_{ijk}^{4}([u],[v],[au+v])\neq\emptyset.$ (5)
This is equivalent to the existence of $\zeta]=[bu+v+cw]$ in $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{u,v}}$ satisfying the following conditions.
  1. C1.
    There exist $\sigma\in U(3,q^{2})$ and $r,s,t\in F^{\times}$ such that
    $\sigma(u,v,b_{i}u+v+c_{i}w)=(r(bu+v+cw),sv,t(au+v)).$ (6)
  2. C2.
    There exist $\tau\in U(3,q^{2})$ and $r_{1},s_{1},t_{1}\in F^{\times}$ such that
    $\tau(u,v,b_{j}u+v+c_{j}w)=(r_{1}u,s_{1}(bu+v+cw),t_{1}(au+v)).$ (7)
Since $\tau\in GL(3,q^{2})$ is in $U(3,q^{2})$ if and only if $B(\tau(x),\tau(y))=B(x,y)$ for $x,y\in\{u,v,w\}$, Condition C2 is equivalent to
$0=\overline{s_{1}}r_{1}-1=s_{1}-t_{1}=c\overline{c}s_{1}+b_{j}r_{1}+bs_{1}-at_{1},$ (8)
for some $r_{1},s_{1},t_{1}\in F^{\times}$. In turn, this is equivalent to
$b_{j}(\overline{b}+a)^{-1}\in F_{0}^{\times}.$ (9)
Indeed, Equation (8) implies $b_{j}(\overline{b}+a)^{-1}\in F_{0}^{\times}$. Conversely, if $b_{j}(\overline{b}+a)^{-1}\in F_{0}^{\times}$, then we may choose $s_{1}\in F^{\times}$ such that $b_{j}=s_{1}\overline{s_{1}}(\overline{b}+a)$. Equation (8) is then satisfied by taking $t_{1}=s_{1}$ and $r_{1}=\overline{s_{1}}^{-1}$.
Similarly, Condition C1 holds if and only if
$b_{i}b(\overline{b}+a)a^{-1}\in F_{0}^{\times}.$ (10)
It remains to show that if $p_{ijk}^{4}\neq 0$, then there are exactly $q+1$ elements $\zeta]$ in the orbit $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{u,v}}$ such that Equations (9) and (10) are true. To see this, fix $\zeta]=[bu+v+cw]\in\Gamma_{ijk}^{4}([u],[v],[au+v]).$ Recall that the elements of $([b_{k}u+v+c_{k}w])^{U(3,q^{2})_{u,v}}$ are $[du+v+ew]$, where $d\in bF_{0}^{\times}$, and $e\in F^{\times}$ satisfies $d+\overline{d}+e\overline{e}=0$. For each of the $q+1$ elements $e\in F^{\times}$ satisfying $e\overline{e}=c\overline{c}$, the element $[bu+v+ew]$ satisfies Equations (9) and (10). This accounts for $q+1$ elements of $\Gamma_{ijk}^{4}([u],[v],[au+v])$. To complete the proof, we show that $du+v+ew]\notin\Gamma_{ijk}^{4}([u],[v],[au+v])$ if $b\neq d$. Indeed, let $d=by$ for some $1\neq y\in F_{0}^{\times}$. If $du+v+ew]\in\Gamma_{ijk}^{4}([u],[v],[au+v])$, then Equation (10) yields $b_{i}d(\overline{d}+a)a^{-1}\in F_{0}^{\times}$. Since $b_{i}b(\overline{b}+a)a^{-1}$ is also in $F_{0}^{\times}$, we obtain
$y^{-1}(\overline{b}+a)(a+\overline{b}y)^{-1}=b_{i}b(\overline{b}+a)a^{-1}(b_{i}d(\overline{d}+a)a^{-1})^{-1}=z,$ (11)
for some $z\in F_{0}^{\times}$. Since $y\neq 1$, Equation (11) implies $yz\neq 1$. Consequently, Equation (11) yields $\frac{a}{\overline{b}}=\frac{1-y^{2}z}{zy-1}=\frac{\overline{a}}{b}=-\frac{a}{b}$. This implies $b+\overline{b}=0$, which is impossible. ∎
As a consequence of Theorem 3.2, we obtain a minimal ternary subalgebra generated by the hypermatrix $A_{4}$ of the AST from $U(3,q^{2})$.
Corollary 3.2.
Let $X=\left\{R_{0},R_{1},R_{2},R_{3},\ldots,R_{q+4}\right\}$ be the AST from the $2$-transitive action of $U(3,q^{2})$ on $\mathbb{P}C$, where the relations are as given in Equations (3) and (4). If $\left\{A_{0},A_{1},A_{2},A_{3},\ldots A_{q+4}\right\}$ is the set of corresponding adjacency hypermatrices of $X$, then $\operatorname{Span}\left\{A_{4}\right\}$ is a ternary subalgebra of $\operatorname{Span}\left\{A_{4},\ldots A_{q+4}\right\}$.
Proof.
By lines $6$ and $9$ in Table 1, we have
$A_{4}A_{4}A_{4}=\sum_{\ell=4}^{q+4}{p_{444}^{\ell}A_{\ell}}=p_{444}^{4}A_{4}+\sum_{\ell=5}^{q+4}{p_{444}^{\ell}A_{\ell}}=(q-2)A_{4}.\qed$
When the characteristic of the underlying field is two, several intersection numbers vanish.
Corollary 3.3.
Let $q$ be a power of two and $X$ be the AST from the $2$-transitive action of $U(3,q^{2})$ on $\mathbb{P}C$. If the relations of $X$ are as given in Equations (3) and (4), then
$p_{k\ell k}^{\ell}=p_{\ell kk}^{\ell}=p_{kk\ell}^{\ell}=p_{\ell\ell\ell}^{\ell}=0,$
whenever $k,\ell\in\{5,\ldots,m\}$ are distinct.
Proof.
We only show that $p_{k\ell k}^{\ell}=0$ if $k,\ell\in\{5,\ldots,m\}$ are distinct. Assume on the contrary that $p_{k\ell k}^{\ell}\neq 0$. Choose any $([u],[v],[bu+v+cw])\in R_{\ell}$ and suppose $([u],[v],[hu+v+ew])\in R_{k}$ satisfies the conditions of Equation $13$ of Table 1. As a consequence, we have
$hhb^{-1}b\in F_{0}.$
It follows that $h^{2}\in F_{0}$. Since $h+\overline{h}+e\overline{e}=0$, we obtain
$h+e\overline{e}$ $=-\overline{h},\text{ and }$
$h$ $=-\overline{h}-e\overline{e}.$
Multiplying the corresponding sides of the above equations yields $h^{2}+he\overline{e}=\overline{h}^{2}+\overline{h}e\overline{e}$. Since $h^{2}=\overline{h}^{2}$, it follows that $h=\overline{h}$. That is, $h\in F_{0}^{\times}$. Since $\operatorname{char}F=2$ and $k\geq 5$, we obtain a contradiction. ∎